# SCMM: Calibrating Cross-modal Representations for Text-Based Person Search

### v7

@@ 4.4 Ablation Studies
4.4.1 Analysis of Model Components
To fully validate the performance of the different components in SCMM, we demonstrate the contributions of each part on CUHK-PEDES, as shown in Table 4. Model 1 and Model 2 show the results using ResNet [42] and ViT [37] as the image encoder, respectively, without the Sew Calibration and MCM losses. We utilize Model 2 as our baseline, and CMPM and CMPC [16] are used as loss functions in the baseline experiment. First, we compare ResNet-50 and ViT-Base as the image encoder and observe a performance improvement of 1.90%, 1.03%, and 0.69% on Rank1, Rank5, and Rank10, respectively. Based on this, we adopt ViT as the image encoder in the following experiments. We then validate the sew calibration and MCM losses compared to the baseline. From Models 2-4, we observe that sew calibration loss brings marginal improvement, and a common distribution can provide a better basis for optimizing the embedding space for fine-grained cross-modal recognition. Moreover, Model 3 represents the sew calibration loss with the fixed margin 0.5, while Model 4 represents the sew calibration loss with adaptive margins. We observe that Model 4 achieves better results than a fixed manual margin, no matter how we tune the margin value, which demonstrates the effectiveness of adaptive constraints. We also find that MCM gives another substantial improvement to the baseline from Model 2 to Model 5, indicating that the text-image detail mining capability is critical. Notably, when the two components are used jointly in Model 6, SCMM continues to improve and outperform the baseline by 4.57%, 2.99%, and 2.37% on the three metrics, respectively. Importantly, the results demonstrate that obtaining well-distributed image-text features in the common embedding space is essential for reducing the cross-modal gap. The consistent improvement in each component of the method demonstrates our effectiveness.
@@ 4.4.1 Analysis of Model Components
4.4.2 Impact of Sew Calibration Loss for Reducing Cross-modal Gap
Benefiting from the sew calibration loss, which reduces the cross-modal gap, the learned representation distributions are closer than with the basic loss. To demonstrate this, we illustrate a comparison of singular value decomposition on cross-modal representations, inspired by [3]. Specifically, Fig. 4(a) and (c) present the baseline and SCMM’s singular value decomposition for the text and image modalities. We computed the distance between these two modalities at specific singular values and reflected them in Fig. 4(b) and (d). Intuitively, the smaller the distance between different modality features, the closer the distribution learned by the model. We find that the sew calibration loss performs better representation distributions in the common embedding space. It ensures closer cross-modal representation distributions and a smaller inter-modal gap than the baseline.
4.4.3 Impact of Masked Caption Modeling
The masked caption modeling operation in fine-grained cross-modal interaction is critical in our framework. It comprises tokens masking, attention module, and masked tokens prediction. As shown in Table 5, we explore the effectiveness of MCM components on the CUHK-PEDES. First, only masking on the input text tokens without a reconstruction task behaves like a random erase text augmentation. Remarkably, the augmentation can already bring +1.36% marginal improvement and reach 63.65% on Rank1. It shows that the details provided by word tokens are helpful for fine-grained recognition.
Next, if we only utilize attention mechanism to enhance cross-modal representations without the mask caption modeling, it can achieve 63.83% on Rank1. We explain this improvement as the cross-attention brings details from sequence tokens to the CLS token. Meanwhile, image features also contribute a lot to this process.
@@ 4.4.3 Impact of Masked Caption Modeling
Table 5: MCM components on the CUHK-PEDES dataset.
@@
4.4.4 Analysis of Generalisability Validation
Table 6: Performance analysis for cross-domain validation.
@@ 4.4.4 Analysis of Generalisability Validation
To better understand the contributions of MCM in our framework, we conduct a domain generalization analysis on the three benchmarks to demonstrate our generalization performance. In Table 6, CUHK $\Rightarrow$ ICFG and CUHK $\Rightarrow$ RSTP indicate using the model trained on the CUHK-PEDES dataset to infer their test sets and vice versa. Compared to the baseline, we can observe a performance improvement in the three metrics for ICFG-PEDES and RSTPReid. The fine-grained features mined by MCM are more resistant to overfitting. The improvement in domain generalization shows the capability of our framework in learning generic cross-modal information, which is essential to solve the fine-grained modal heterogeneity.
4.4.5 Effect of Margin Parameters
In Sec.  3.1, we introduce the adaptive margins to constrain cross-modal learning using Sew Calibration loss with constraints. We find that compared to the manually fixed parameters (SEW-f), our adaptive margins (SEW-a) produce better results as shown in Fig. 5, where we investigate the effect of manual margin in different settings on the CUHK-PEDES [8] on Rank1 accuracy. The best performance of 65.41% is achieved when $\mathcal{M}=0.5$, while the worst model performance of 62.29% is observed when $\mathcal{M}=0$. We note that the performance of the model first improves as the margin increases, which is due to our bi-directional margin compressing the same pedestrian target compactly. However, we find that the model performance does not continue to improve as the margin increases, and we observe an inflection point around margin 0.5. When the margin is too large, the tight constraints make the model too hard to train.
Therefore, it is critical to utilize quality-guided adaptive margins according to the textual information. In this regard, we set the upper and lower margin bounds ($\mathcal{M}_{min}=0.4$ and $\mathcal{M}_{max}=0.6$). In this range, our Sew Calibration loss with adaptive constraints obtains the best performance, which a manually fixed margin cannot achieve, regardless of how the margin value is tuned.
4.4.6 Effect of Mask Ratio
Inspired by Masked Language Modeling [36], we propose the MCM loss which leverages a masked captions prediction task to establish detailed and generic relationships between textual and visual parts. Our method requires setting the mask ratio in the MCM. To investigate the effect of the mask ratio on model performance and generalization learning ability, we conducted ablation experiments, and the results are shown in Table 7. When the mask ratio is set to 0, which is equivalent to the baseline, the performance is the worst. Compared to the baseline, our method with a 0.1 mask ratio achieves the best results, with scores of 65.16%, 81.95%, and 87.93% on Rank1, Rank5, and Rank10, respectively.
Table 7: Ablation results of mask ratio on the CUHK-PEDES.
@@ 4.4.6 Effect of Mask Ratio
4.5 Qualitative Analysis
4.5.1 Visualization of Attention Map
@@
We visualize attention maps on the CUHK-PEDES test dataset to demonstrate model capability in learning image-text correspondence. As shown in Fig. 6, compared to the baseline, we can observe that the visualization results obtained by SCMM are more apparent and refined. We conduct the word-level visualization to validate further the ability to perform fine-grained interaction. We select several keywords in the caption description, e.g., bag, shoes as items, colors, and clothing as modifiers. For example, the visualization results of IDs 8491 and 10648 do not focus on useful detailed information. The key messages in the two images are a man riding bike with a backpack and a woman wearing a sleeveless white dress. We can observe that SCMM successfully captures the key detail information compared to the baseline. From the word-level results, we can also observe that SCMM learns the critical parts in the cross-modal correspondence.
Figure 7: Presentation of cross-modal representation distributions on CUHK-PEDES test dataset. Different colors correspond to the different target ID. Best viewed in color.
Figure 8: Top 5 retrieval results on CUHK-PEDES. Matched target person images are marked with red rectangles. and subcaptions show the text queries used for retrieval.
4.5.2 Visualization of Image-text Feature Distribution
@@
4.5.3 Visualization of Top 5 Search Qualitative Results
To evaluate the performance of our method, we randomly selected several sets of search results from our method and visualized them in Fig. 8. For each target image, we show the top 5 most similar search results, with the similarity scores after re-ranking processing displayed above each image. Correct search results are indicated by red rectangles, and the number in the last column shows how many pedestrian targets with the same ID have been recalled. Taking ID 5219 for an example, our method correctly retrieved all three target images, with similarity scores of 1.2563, 1.2100, and 0.9775, respectively. For the last two unmatched images, our method gave similarity scores of 0.5174 and 0.5011, respectively. Although these images have some similarities with the positive samples, they do not match the text description in other details, and our method learned to differentiate them. The last row in Fig. 8 shows a hard case, where the man in the query image has extremely similar attributes to several other pedestrians. Despite, our method is able to correctly rank 2/3 of the right person images as the top 2 search results.
5 Conclusion
In this work, we introduce SCMM, a novel AI framework for text-based person search that addresses fundamental challenges in cross-modal learning. Our method introduces two innovative components: sew calibration loss with adaptive margins and masked caption modeling loss that establishes fine-grained correspondences through transformer-based architectures. These complementary strategies bridge the inter-modal gap while capturing detailed relationships essential for person identification. Extensive experimental validation demonstrates SCMM’s superior performance, achieving state-of-the-art results of 73.81%, 64.25%, and 57.35% Rank1 accuracy on CUHK-PEDES, ICFG-PEDES, and RSTPReID respectively. SCMM’s streamlined dual-encoder architecture prioritizes computational efficiency, making it suitable for real-time applications. Future work will explore incorporating additional modalities and developing more sophisticated adaptive mechanisms for multimodal learning in person re-identification tasks.
CRediT authorship contribution statement
Jing Liu: Conceptualization, Methodology, Formal analysis, Writing – original draft, Writing – review & editing, Visualization. Donglai Wei: Methodology, Software, Validation, Formal analysis, Investigation, Writing – review, Project administration & editing. Yang Liu: Conceptualization, Methodology, Investigation, Writing – review & editing, Visualization. Sipeng Zhang: Writing – review & editing, Funding acquisition. Tong Yang: Writing – review & editing, Funding acquisition. Wei Zhou: Conceptualization, Writing – review & editing. Weiping Ding: Methodology, Writing – review & editing. Victor C. M. Leung: Supervision, Writing – review & editing, Project administration.
Data availability
Data will be made available on request.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used generative AI and AI-assisted technologies for writing enhancement and language polishing to improve the clarity and readability of the manuscript. After using these tools, the authors have reviewed and edited the content as needed and take full responsibility for the content of the publication.
Acknowledgements
This work was supported in part by the Mitacs Elevate under Grant No. IT44479; the Guangdong Pearl River Talents Recruitment Program under Grant No. 2019ZT08X603; and the Guangdong Pearl Rivers Talent Plan under Grant No. 2019JC01X235. We sincerely thank all the editors and anonymous reviewers for their careful work and thoughtful suggestions that have helped improve this paper substantially.

### v8

@@ IV-D Ablation Studies
Notes: Res and ViT denote ResNet and Vision Transformer as image encoders, respectively. Sew-F and Sew-A refer to the fixed and adaptive margin versions of the sew calibration loss, respectively. Model 2 is our baseline.
IV-D1 Analysis of Model Components
To validate SCMM’s components, we evaluate their contributions on CUHK-PEDES (Table IV). Models 1 and 2 compare ResNet-50 and ViT-Base encoders, with ViT showing superior performance (+1.90% Rank1). Using Model 2 as the baseline, Models 3 and 4 demonstrate that the sew calibration loss improves alignment, with adaptive margins (Model 4) outperforming fixed margins (Model 3). Model 5 shows that MCM significantly boosts performance by mining fine-grained details. Notably, the full SCMM framework (Model 6) combines these components to achieve the best results, outperforming the baseline by 4.57% in Rank1 accuracy. The consistent improvement confirms that jointly optimizing global alignment and fine-grained correspondence is essential for effective cross-modal representation learning.
@@ IV-D1 Analysis of Model Components
IV-D2 Impact of Sew Calibration Loss for Reducing Cross-modal Gap
Benefiting from the sew calibration loss, which reduces the cross-modal gap, the learned representation distributions are closer than with the basic loss. To demonstrate this, we illustrate a comparison of singular value decomposition on cross-modal representations [xie2025multiscale]. Specifically, Fig. 4(a) and (c) present the baseline and SCMM’s singular value decomposition for the text and image modalities. We computed the distance between these two modalities at specific singular values and reflected them in Fig. 4(b) and (d). Intuitively, the smaller the distance between different modality features, the closer the distribution learned by the model. We find that the sew calibration loss performs better representation distributions in the common embedding space. It ensures closer cross-modal representation distributions and a smaller inter-modal gap than the baseline.
IV-D3 Impact of Masked Caption Modeling
The masked caption modeling operation in fine-grained cross-modal interaction is critical in our framework. It comprises token masking, an attention module, and masked tokens prediction. As shown in Table V, we explore the effectiveness of MCM components on the CUHK-PEDES. First, only masking on the input text tokens without a reconstruction task behaves like a random erase text augmentation. Remarkably, the augmentation can already bring +1.36% marginal improvement and reach 63.65% on Rank1. It shows that the details provided by word tokens are helpful for fine-grained recognition. Next, if we only utilize the attention mechanism to enhance cross-modal representations without the mask caption modeling, it can achieve 63.83% on Rank1. We explain this improvement as the cross-attention brings details from sequence tokens to the CLS token. Meanwhile, image features also contribute a lot to this process.
@@ IV-D3 Impact of Masked Caption Modeling
TABLE V: MCM components on the CUHK-PEDES dataset.
@@
IV-D4 Analysis of Generalisability Validation
TABLE VI: Performance analysis for cross-domain validation.
@@ IV-D4 Analysis of Generalisability Validation
To better understand the contributions of MCM in our framework, we conduct a domain generalization analysis on the three benchmarks to demonstrate our generalization performance. In Table VI, CUHK $\Rightarrow$ ICFG and CUHK $\Rightarrow$ RSTP indicate using the model trained on the CUHK-PEDES dataset to infer their test sets and vice versa. Compared to the baseline, we can observe a performance improvement in the three metrics for ICFG-PEDES and RSTPReid. The fine-grained features mined by MCM are more resistant to overfitting. The improvement in domain generalization shows the capability of our framework in learning generic cross-modal information, which is essential to solving the fine-grained modal heterogeneity.
IV-D5 Effect of Margin Parameters
In Sec.  III-A, we introduce the adaptive margins to constrain cross-modal learning using Sew Calibration loss with constraints. We find that compared to the manually fixed parameters (SEW-f), our adaptive margins (SEW-a) produce better results, as shown in Fig. 5, where we investigate the effect of manual margin in different settings on the CUHK-PEDES [gnarnn] on Rank1 accuracy. The best performance of 65.41% is achieved when $\mathcal{M}=0.5$, while the worst model performance of 62.29% is observed when $\mathcal{M}=0$. We note that the performance first improves as the margin increases, which is due to our bi-directional margin compressing the same pedestrian target compactly. However, we find that the performance does not continue to improve as the margin increases, and we observe an inflection point around a margin of 0.5. When the margin is too large, the tight constraints make the model too hard to train. Therefore, it is critical to utilize quality-guided adaptive margins according to the textual information. In this regard, we set the upper and lower margin bounds ($\mathcal{M}_{min}=0.4$ and $\mathcal{M}_{max}=0.6$). In this range, our Sew Calibration loss with adaptive constraints obtains the best performance, which a manually fixed margin cannot achieve, regardless of how the margin value is tuned.
IV-D6 Effect of Mask Ratio
Inspired by Masked Language Modeling [lee2023lanobert], we propose the MCM loss which leverages a masked captions prediction task to establish detailed and generic relationships between textual and visual parts. Our method requires setting the mask ratio in the MCM. To investigate the effect of the mask ratio on model performance and generalization learning ability, we conducted ablation experiments, and the results are shown in Table VII. When the mask ratio is set to 0, which is equivalent to the baseline, the performance is the worst. Compared to the baseline, our method with a 0.1 mask ratio achieves the best results, with scores of 65.16%, 81.95%, and 87.93% on Rank1, Rank5, and Rank10, respectively.
TABLE VII: Ablation results of mask ratio on the CUHK-PEDES.
@@ IV-D6 Effect of Mask Ratio
IV-E Qualitative Analysis
IV-E1 Visualization of Attention Map
@@
We visualize attention maps on the CUHK-PEDES test dataset to demonstrate model capability in learning image-text correspondence. As shown in Fig. 6, compared to the baseline, we can observe that the visualization obtained by SCMM is more apparent and refined. We conduct the word-level visualization to validate further the ability to perform fine-grained interaction. We select several keywords in the caption description, e.g., bag, shoes, colors, and clothing as modifiers. For example, the visualization of IDs 8491 and 10648 does not focus on useful detailed information. The key messages in the two images are a man riding a bike with a backpack and a woman wearing a sleeveless white dress. We can observe that SCMM successfully captures the key detail information compared to the baseline. From the word-level results, we observe that SCMM learns the critical parts in the cross-modal correspondence.
Figure 7: Presentation of cross-modal representation distributions on CUHK-PEDES test dataset. Different colors correspond to the different target ID.
IV-E2 Visualization of Image-text Feature Distribution
@@
V Conclusion
We proposed SCMM, a novel framework for text-based person search that calibrates cross-modal representations. SCMM integrates a sew calibration loss with quality-guided adaptive margins for global alignment and a masked caption modeling loss for fine-grained correspondence. The dual-encoder architecture with a training-only decoder balances efficiency and performance. Extensive experiments on three benchmarks demonstrate SCMM achieves state-of-the-art results (e.g., 73.81% Rank1 on CUHK-PEDES). Ablation studies confirm the synergistic effects of our proposed components in bridging the inter-modal gap and capturing detailed attributes. Future work will explore incorporating additional modalities and developing more sophisticated adaptive learning mechanisms.